OpenStudy (anonymous):

-2/3y+12=-8 its negative 2 over 3

OpenStudy (anonymous):

-2/3y=-8-12 -2/3y=-20 -2y=-20*3 -2y=-60 y=-60/(-2) y=30

OpenStudy (anonymous):

ohhh Thank You. Can you help me with another please?